# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix A Background: Distribution Aware General Wasserstein Mechanism
$\hat{\epsilon}=\frac{d\log\frac{\beta_{k,\lambda}\pi}{\lambda}\overline{Q}_{kq(\alpha-1)/2}\left(1+\left(\frac{\Delta^{dkq(\alpha-1)}_{G}}{d\lambda}\right)^{2}\right)}{q(\alpha-1)}$
with $\Delta^{dkq(\alpha-1)}_{G}$ given by (24).
The DAGWM is proved to offer more utility than the GWM at no additional privacy cost [pierquin2024renyi].
Curse of dimensionality for DAGWM.
Although DAGWM replaces $\infty$-Wasserstein sensitivities by $p$-Wasserstein distances in (23), it does not eliminate the curse of dimensionality inherent to optimal transport. For a query $f:\mathcal{X}\to\mathbb{R}^{d}$, the calibration term $\Delta^{\zeta,q,\alpha}_{G}$ in (23) and $\Delta^{dkq(\alpha-1)}_{G}$ in (24) depend on high-dimensional $W_{p}$-distances, $W_{p}\big(P(f(X)\mid s_{i},\theta),P(f(X)\mid s_{j},\theta)\big)$ in $\mathbb{R}^{d}$. In practice, these quantities must be estimated from empirical distributions based on $n$ samples. Classical results on empirical Wasserstein distances imply that, for probability measures on $\mathbb{R}^{d}$ with bounded support and finite $p$-th moments, the expected estimation error obeys
$\mathbb{E}\big[\,\big|W_{p}(\hat{\mu}_{n},\mu)-W_{p}(\mu,\nu)\big|\,\big]\;\gtrsim\;c_{d}\,n^{-1/d},$
for some dimension-dependent constant $c_{d}>0$. Thus, to achieve a target accuracy $\varepsilon$ in the Wasserstein sensitivity (and hence in the DAGWM calibration), the required sample size scales at least as $n=\Omega(\varepsilon^{-d})$. Moreover, computing $W_{p}$ (or a sufficiently accurate approximation) between two $n$-point empirical measures in $\mathbb{R}^{d}$ entails solving a high-dimensional optimal transport problem, whose complexity also grows rapidly with both $n$ and $d$. Consequently, while DAGWM improves the distribution-awareness of RPP, it still inherits the statistical and computational curse of dimensionality of high-dimensional Wasserstein distances, and may become impractical when $f(X)$ is moderately or highly dimensional.
Appendix B Average vs. Joint SRPP
@@
$\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}):=\int_{\mathbb{S}^{d-1}}\mathtt{D}_{\alpha}\big(\pi_{u}{}_{\#}P\,\big\|\,\pi_{u}{}_{\#}Q\big)d\omega(u),$
@@
$)].\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}):=\frac{1}{\alpha-1}\log\mathbb{E}_{U\sim\omega}\Big[\exp\big((\alpha-1)\mathtt{D}_{\alpha}\big(\pi_{U}{}_{\#}\mathrm{P}\big\|\pi_{U}{}_{\#}\mathrm{Q}\big$$\big)\Big$].
@@
$\mathtt{Slice}(M):x\to(U,\pi_{U}(\mathcal{M}(x))),\quad U\sim\omega,$
then the Joint-SRD coincides with the standard Rényi divergence
$\mathtt{JSD}^{\nu}_{\alpha}\!\big(\mathrm{P}\|\mathrm{Q}\big)=\mathtt{D}_{\alpha}\big(\widetilde{\mathrm{P}}(U,\pi_{U}(Z_{\mathrm{P}}))\,\big\|\,\widetilde{\mathrm{P}}(U,\pi_{U}(Z_{\mathrm{Q}}))\big),$
where $Z_{\mathrm{P}}\sim\mathrm{P}$ and $Z_{\mathrm{Q}}\sim\mathrm{Q}$. In this sense, Joint-SRPP is exactly RPP applied to the augmented, “sliced” observation $(U,\pi_{U}(\mathcal{M}(X)))$, whereas Ave-SRPP only controls the mean per-direction divergence.
@@
and the same inequalities hold for the corresponding output laws under any fixed Pufferfish model $(\mathcal{S},\mathcal{Q},\Theta)$ ((Proposition 4.1)). Thus, at the level of numerical certificates, Joint-SRPP is always at least as “conservative" as Ave-SRPP for the same $(\alpha,\epsilon,\omega)$, and both are upper-bounded by the unsliced RPP divergence.
Why do we introduce both notions?
@@
Although Ave-SRPP and Joint-SRPP are closely related, they serve complementary purposes:
@@
  • Ave-SRPP for geometry-aware average protection. Ave-SRPP directly reflects the geometry induced by the slicing distribution $\omega$ and is naturally aligned with sliced Wasserstein mechanism privatization, where one averages per-direction sensitivities. Formally, it controls the expected sliced Rényi divergence when the projection direction is drawn from $\omega$. Equivalently, it quantifies privacy risk as if one only observed a single one-dimensional view sampled from $\omega$, with average Rényi cost at most $\epsilon$. This “single-slice observer" is purely an analytic device: the actual mechanism always outputs the full $d$-dimensional vector, and the sliced view is used only to define a geometry-aware privacy budget. In practice, this makes Ave-SRPP a natural calibration target, often yielding tighter, more utility-friendly noise levels than enforcing a worst-case bound simultaneously in every direction.

### v3

@@ A.5.3 Limitations of Sinkhorn-Wasserstein Mechanism
Practical implication.
[…]
Appendix B Properties of SRPP
In this section, we provide detailed discussions of Ave-SRPP and Joint-SRPP. Ave-SRPP and Joint-SRPP use the same ingredients (per-slice Rényi divergences along 1-D projections) but aggregate them in different ways, leading to different strengths. Ave-SRPP averages the per-slice Rényi divergences over directions $u\sim\omega$, so it controls the mean distinguishability of secrets across slices (good for mean geometry across slices). Joint-SRPP instead can be interpreted as bounding the Rényi divergence of the joint observation of slice index $V\sim\omega$ and projection $\Psi^{V}_{\#}(\mathcal{M}(X))$, which captures and penalizes rare, high-risk directions that Ave-SRPP alone can hide in the average. Joint-SRPP is generally a stronger requirement than Ave-SRPP: satisfying Joint-SRPP implies the corresponding averaged bound, but not conversely.
Sec. B.1 shows the ordering of Ave-SRD, Joint-SRD, and the standard Rényi divergence. In Sec. B.2, we characterize the post-processing immunity property of SRPP notions. Sec. B.3 further compares Ave-SRPP and Joint-SRPP. Finally, we show that our SRPP frameworks imply their corresponding sliced Pufferfish Privacy frameworks in Sec. B.4.
B.1 Ordering
Our two SRPP frameworks sit between each other and the standard RPP.
Proposition B.1.
$\mathtt{AveSD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{JSD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{D}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big).$
The ordering in Proposition B.1 only compares the numerical tightness of upper bounds and does not induce a straightforward ranking of privacy strength between the three notions. For example, an $(\alpha,\epsilon)$-RPP guarantee implies the corresponding $(\alpha,\epsilon,\omega)$-Joint/Ave-SRPP guarantees by Proposition B.1, but the converse need not hold. Moreover, different choices of $\omega$ and different calibration strategies for the noise can lead to mechanisms with the same bound on $\mathtt{AveSD}^{\omega}_{\alpha}$ or $\mathtt{JSD}^{\omega}_{\alpha}$ but very different protections at the level of secrets and priors. For this reason, the inequalities in Proposition B.1 should be interpreted as relations between divergence-based certificates, rather than as an absolute ranking of which framework is "more private".
Proof of Proposition B.1.
$D(u):=\mathtt{D}_{\alpha}\bigl(\Psi^{u_{\ell}}_{\#}\mathrm{P}\|\Psi^{u_{\ell}}_{\#}\mathrm{Q}\bigr),\qquad\ell=1,\dots,m.$
By Definition 4.1, the average sliced Rényi divergence is
$\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})=\int_{\mathbb{S}^{d-1}}D(u)\ d\omega(u)$
By Definition 4.3, the joint sliced Rényi divergence is
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})=\frac{1}{\alpha-1}\log\Bigl(\int_{\mathbb{S}^{d-1}}\exp\bigl((\alpha-1)D(u)\ d\omega(u)\bigr)\Bigr).$
Part 1: $\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\leq\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$.
Let $Z$ be the random variable $Z=D(U)$ when $U\sim\omega$. Then,
$\mathbb{E}[Z]=\int_{\mathbb{S}^{d-1}}D(u)\ d\omega(u)=\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}),$
and
Thus, by taking log and dividing by $(\alpha-1)$, we get
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\geq\mathbb{E}[Z]=\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}).$
Define the randomized post-processing channel $\mathsf{Slice}$ as follows: given $Y\in\mathbb{R}^{d}$, draw $U\sim\omega$ independently of $Y$, then output $\mathsf{Slice}(Y):=(U,\langle Y,U\rangle)$. Let $\mathrm{P}^{\mathsf{Slice}}$ and $\mathrm{Q}^{\mathsf{Slice}}$ denote the laws of $\mathsf{Slice}(Y)$ when $Y\sim\mathrm{P}$ and $Y\sim\mathrm{Q}$, respectively.
For each direction $u$, the conditional law of $\langle Y,u\rangle$ under $Y\sim\mathrm{P}$ is $\Psi^{u}_{\#}\mathrm{P}$, and similarly under $\mathrm{Q}$ it is $\Psi^{u}_{\#}\mathrm{Q}$. By the definition of Rényi divergence for product/mixture measures with the same mixing law $\omega$, we have
$\mathtt{D}_{\alpha}(\mathrm{P}^{\mathsf{Slice}}\|\mathrm{Q}^{\mathsf{Slice}})$
$=\frac{1}{\alpha-1}\log\!\int_{\mathbb{S}^{d-1}}\exp\!\Big((\alpha-1)\mathtt{D}_{\alpha}(\Psi^{u}_{\#}\mathrm{P}\,\|\,\Psi^{u}_{\#}\mathrm{Q})\Big)\,d\omega(u)$
$=\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}).$
Since $\mathsf{Slice}$ is a (data-independent) post-processing channel applied to $Y$, the data-processing inequality for Rényi divergence (an $f$-divergence for $\alpha>1$) yields
$\mathtt{D}_{\alpha}(\mathrm{P}^{\mathsf{Slice}}\|\mathrm{Q}^{\mathsf{Slice}})\leq\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q}),$
and hence $\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})\leq\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q})$.
By combining Part 1 and Part 2, we can claim
$\mathtt{AveSD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{JSD}^{\omega}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big)\leq\mathtt{D}_{\alpha}\big(\mathcal{M}^{\theta}_{s_{i}}\|\mathcal{M}^{\theta}_{s_{j}}\big),$
B.2 Post-Processing Immunity
A fundamental closure property of DP and RPP is post-processing immunity (PPI): the privacy loss of a mechanism cannot be increased by arbitrary data-independent post-processing [pierquin2024renyi, Dwork2014]. Formally, for any Markov kernel (randomized mapping) $K$ that is independent of the dataset, post-processing acts on the endpoint distributions of a release. Since the Rényi divergence $\mathtt{D}_{\alpha}(\cdot\|\cdot)$ satisfies the data-processing inequality (DPI),
$\mathtt{D}_{\alpha}(PK\,\|\,QK)\ \leq\ \mathtt{D}_{\alpha}(P\,\|\,Q),$
any privacy certificate expressed in terms of (upper bounds on) $\mathtt{D}_{\alpha}$ is automatically closed under post-processing.
Post-processing immunity has two aspects. (i) True privacy loss never increases: for any data-independent $K$, $\mathtt{D}_{\alpha}(PK\,\|\,QK)\leq\mathtt{D}_{\alpha}(P\,\|\,Q)$ by DPI. (ii) Safe certificate reuse for a downstream pipeline: one may upper-bound the privacy loss of a post-processing pipeline using a certificate for the initial mechanism. For SRPP, this reuse must be interpreted through the relationship between sliced divergences and the full-dimensional Rényi divergence ( Prop. B.1): SRPP/SRD is a quantification language for the full release, and changing the representation along the pipeline can change the numerical SRPP value even though the underlying distinguishability (and hence privacy loss in the sense of (i)) cannot increase. Consequently, reusing SRPP parameters for an end-to-end pipeline can be valid but may be overly conservative when the quantification geometry matters.
In privacy, PPI should be understood primarily as a semantic closure property: for any data-independent post-processing $K$ applied to the true release, the post-processed pipeline cannot be more distinguishable than the original release in the underlying divergence sense (DPI). This is stronger than (and logically prior to) any particular parameterization rule such as "reusing the same numeric privacy parameters", which is merely a convenient upper-bounding convention and can be very conservative (or not directly meaningful) when the quantification geometry changes (as in SRPP).
SRPP’s post-processing discussion is ultimately rooted in the data-processing inequality (DPI) for the standard Rényi divergence $\mathtt{D}_{\alpha}$. Concretely, SRPP quantifies privacy by applying a fixed (profile-determined) slicing/observation channel to the endpoint distribution of the (full-dimensional) released output, and then evaluating $\mathtt{D}_{\alpha}$ on the induced sliced laws: Ave-SRPP takes the $\omega$-average of the resulting per-direction divergences, while Joint-SRPP applies a log-moment aggregation that is equivalent to a single $\mathtt{D}_{\alpha}$ on an augmented observation. Accordingly, for any data-independent post-processing kernel $K$ applied to the true release $Y=\mathcal{M}(X)$, one first obtains the post-processed endpoint laws $PK$ and $QK$, and SRPP/SRD is then evaluated on $PK$ versus $QK$ under the same slice profile. Any nonconservation of the SRPP numerical value under post-processing reflects this change in endpoint distributions (and the geometry induced by the chosen profile), rather than a failure of post-processing immunity: the underlying distinguishability cannot increase due to DPI.
Let $P=\Pr[\mathcal{M}(X)\mid s_{i},\theta]$ and $Q=\Pr[\mathcal{M}(X)\mid s_{j},\theta]$ be the endpoint distributions on $\mathbb{R}^{d}$, and let $K$ be any data-independent post-processing kernel on $\mathbb{R}^{d}$. Then DPI gives $\mathtt{D}_{\alpha}(PK\|QK)\leq\mathtt{D}_{\alpha}(P\|Q)$ (true privacy loss never increases). Moreover, by Prop. B.1,
$\mathtt{AveSD}^{\omega}_{\alpha}(PK\|QK)\leq\mathtt{JSD}^{\omega}_{\alpha}(PK\|QK)\leq\mathtt{D}_{\alpha}(PK\|QK),$
so any full-dimensional Rényi/RPP certificate for $\mathcal{M}$ remains a valid upper bound on the SRPP/SRD risk of the post-processed pipeline $K\circ\mathcal{M}$. In general, however, the SRPP values $\mathtt{AveSD}^{\omega}_{\alpha}(\cdot\|\cdot)$ and $\mathtt{JSD}^{\omega}_{\alpha}(\cdot\|\cdot)$ need not be numerically conserved under arbitrary $K$, because they quantify endpoint distinguishability through a fixed slice profile.
B.3 Average vs. Joint SRPP
@@ B.2 Post-Processing Immunity
$\mathtt{AveSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q}):=\int_{\mathbb{S}^{d-1}}\mathtt{D}_{\alpha}\big(\Psi^{u}_{\#}P\,\big\|\,\Psi^{u}_{\#}Q\big)d\omega(u),$
@@
$\mathtt{JSD}^{\omega}_{\alpha}(\mathrm{P}\|\mathrm{Q})$
$=\frac{1}{\alpha-1}\log\mathbb{E}_{U\sim\omega}\Big[\exp\big((\alpha-1)\mathtt{D}_{\alpha}\big(\Psi^{U}_{\#}\mathrm{P}\big\|\Psi^{U}_{\#}\mathrm{Q}\big)\big)\Big].$
@@
In particular, define the randomized sliced observation channel $\mathsf{Slice}:\mathbb{R}^{d}\to\mathcal{U}\times\mathbb{R}$ by
$\mathsf{Slice}(y):=(U,\Psi^{U}(y)),\qquad U\sim\omega\ \text{independent of }y.$
Let $\widetilde{\mathrm{P}}:=\mathsf{Slice}_{\#}\mathrm{P}$ and $\widetilde{\mathrm{Q}}:=\mathsf{Slice}_{\#}\mathrm{Q}$ denote the endpoint laws of $(U,\Psi^{U}(Z_{\mathrm{P}}))$ and $(U,\Psi^{U}(Z_{\mathrm{Q}}))$ when $Z_{\mathrm{P}}\sim\mathrm{P}$ and $Z_{\mathrm{Q}}\sim\mathrm{Q}$, respectively. Then
[…]
In this sense, Joint-SRPP is exactly RPP applied to the augmented observation $(U,\Psi^{U}(\mathcal{M}(X)))$, whereas Ave-SRPP only controls the mean per-direction divergence.
@@
and the same inequalities hold for the corresponding output distributions under any fixed Pufferfish model $(\mathcal{S},\mathcal{Q},\Theta)$ ((Proposition B.1)). Thus, at the level of numerical certificates, Joint-SRPP is always at least as "conservative" as Ave-SRPP for the same $(\alpha,\epsilon,\omega)$, and both are upper-bounded by the unsliced RPP divergence.
@@
Although Ave-SRPP and Joint-SRPP are closely related, they serve complementary purposes.
Ave-SRPP for Geometry-Aware Average Protection.
@@
Ave-SRPP directly reflects the geometry induced by the slicing distribution $\omega$ and is naturally aligned with sliced Wasserstein mechanism privatization, where one averages per-direction sensitivities. Formally, it controls the expected sliced Rényi divergence when the projection direction is drawn from $\omega$. Equivalently, it upper-bounds the expected per-direction divergence for a direction drawn from $\omega$; this is an analytic interpretation only (the mechanism still releases the full $d$-dimensional output). This "single-slice observer" is purely an analytic device: the actual mechanism always outputs the full $d$-dimensional vector, and the sliced view is used only to define a geometry-aware privacy budget. In practice, this makes Ave-SRPP a natural calibration target, often yielding tighter, more utility-friendly noise levels than enforcing a worst-case bound simultaneously in every direction.
Joint-SRPP for Robust, Slice-Aware Guarantees.